# Solving the Optimal Experiment Design Problem with Mixed-Integer Convex Methods

### v5

@@ 0.A.2 Generalized Self-Concordance
$\frac{\operatorname{Tr}\left(V^{-\frac{p+1}{2}}UV^{-1}UV^{-\frac{p+1}{2}}\right)^{\nu/2}}{\operatorname{Tr}\left(V^{-\frac{p+1}{2}}UV^{-1}UV^{-1}UV^{-\frac{p+1}{2}}\right)}$ $=\frac{\operatorname{Tr}\left(LL^{\intercal}\right)^{\nu/2}}{\operatorname{Tr}\left(LHL^{\intercal}\right)}.$
Using Cauchy-Schwartz, we find:
$\operatorname{Tr}\left(LHL^{\intercal}\right)$ $=\left\langle L^{\intercal}L,H\right\rangle$
$\leq\operatorname{Tr}\left(LL^{\intercal}\right)\sqrt{\operatorname{Tr}\left(H^{2}\right)}$
$\frac{\operatorname{Tr}\left(LL^{\intercal}\right)^{\nu/2}}{\operatorname{Tr}\left(LHL^{\intercal}\right)}$ $\geq\frac{\operatorname{Tr}\left(LL^{\intercal}\right)^{\nu/2}}{\operatorname{Tr}\left(LL^{\intercal}\right)\sqrt{\operatorname{Tr}\left(H^{2}\right)}}$
$=\operatorname{Tr}\left(V^{p/2}LL^{\intercal}V^{p/2}\right)$
$\geq\operatorname{Tr}\left(LL^{\intercal}\right)^{\nu/2-3/2}\frac{1}{\sqrt{\left\|V^{p}\right\|_{2}}\sqrt[4]{n}}.$
By the assumption that V is a feasible point, we have $\left\|V\right\|_{2}=\lambda_{\max}(V)\leq a$, thus:
$\frac{\operatorname{Tr}\left(LL^{\intercal}\right)^{\nu/2}}{\operatorname{Tr}\left(LL^{\intercal}\right)\sqrt{\operatorname{Tr}\left(H^{2}\right)}}$ $\geq\operatorname{Tr}\left(LL^{\intercal}\right)^{\nu/2-3/2}\frac{1}{\sqrt{a^{p}}\sqrt[4]{n}}.$
Then,
$\frac{M_{f}(p(p+1))^{\nu/2}}{p(p+1)(p+2)}\frac{\operatorname{Tr}\left(V^{-p}UV^{-1}UV^{-1}\right)^{\nu-2}}{\operatorname{Tr}\left(V^{-p}UV^{-1}UV^{-1}UV^{-1}\right)}$ $\geq\frac{M_{f}(p(p+1))^{\nu/2}}{p(p+1)(p+2)}\operatorname{Tr}\left(LL^{\intercal}\right)^{\nu/2-3/2}\frac{1}{\sqrt{a^{p}}\sqrt[4]{n}}.$
We can select $\nu=3$ and re-express the GSC condition as
11 11 $\leq\frac{M_{f}(p(p+1))^{1/2}}{(p+2)}\frac{1}{\sqrt[4]{a^{2p}n}}.$
$M_{f}$ $\geq\frac{(p+2)\sqrt[4]{a^{2p}n}}{\sqrt{p(p+1)}}.$
Thus, $g$ is $\left(3,\frac{(p+2)\sqrt[4]{a^{2p}n}}{\sqrt{p(p+1)}}\right)$-generalized self-concordant on the set $\left\{X\in\mathbb{S}_{++}^{n}\mid 0\prec X\preccurlyeq aI\right\}$. ∎
Corollary 2
The function $g(X)=\operatorname{Tr}\left(X^{-1}\right)$ is $\left(3,\frac{3\sqrt[4]{a^{2}n}}{\sqrt{2}}\right)$-generalized self-concordant on $\mathrm{dom}(g)=\left\{X\in\mathbb{S}_{++}^{n}\mid 0\prec X\preccurlyeq aI\right\}$.
To ensure that Theorem 3.1 is applicable for the objective of (A-Opt), we show that the maximum eigenvalue of the information matrix is bounded.
Lemma 4
Let $x\in\mathcal{P}$. Then,
$\lambda_{\max}(X(\mathbf{x}))\leq m\max_{j\in[m]}u_{j}\max_{i\in[m]}\left\|\mathbf{v}_{i}\right\|_{2}^{2}.$
where $\mathbf{u}$ denotes the upper bounds and $\mathbf{v}_{i}$ are the rows of the experiment matrix $A$.
Proof
$\lambda_{\max}(X(\mathbf{x}))$ $=\lambda_{\max}\left(\sum\limits_{i=1}^{m}x_{i}\mathbf{v}_{i}\mathbf{v}_{i}^{\intercal}\right).$
By the Courant Fischer Min-Max Theorem, we have:
$\lambda_{\max}(X(\mathbf{x}))$ $\leq\sum\limits_{i=1}^{m}\lambda_{\max}\left(x_{i}\mathbf{v}_{i}\mathbf{v}_{i}^{\intercal}\right).$
Using $\mathbf{x}\leq\mathbf{u}$, yields:
$\lambda_{\max}(X(\mathbf{x}))$ $\leq m\max_{j\in[m]}u_{j}\max_{i\in[m]}\left\|\mathbf{v}_{i}\right\|_{2}^{2}.$
Combing Theorem 3.1 and Lemma 4, yields:
Corollary 3
The function $g(\mathbf{x})=\operatorname{Tr}\left(X(\mathbf{x})^{-p}\right)$ is self-concordant on the convex feasible region $\mathcal{P}$.
Appendix 0.B Notes on Custom Branch-and-Bound Algorithm for the Exact Optimal Design Problem
In this section, we showcase the changes we made to the step-size in Ahipaşaoğlu (2021). As defined in Ahipaşaoğlu (2021), we have
$\omega_{j}$ $=\mathbf{v}_{j}^{\intercal}X(\mathbf{w})^{-1}\mathbf{v}_{j}$
$\zeta_{j}$ $=\mathbf{v}_{j}^{\intercal}X(\mathbf{w})^{-2}\mathbf{v}_{j}$
$\omega_{jk}$ $=\mathbf{v}_{j}^{\intercal}X(\mathbf{w})^{-1}\mathbf{v}_{k}$
$\zeta_{jk}$ $=\mathbf{v}_{j}^{\intercal}X(\mathbf{w})^{-2}\mathbf{v}_{k}.$
D-criterion
In the case of the D-criterion, the update of the objective function results in
$\log\det(X(\mathbf{w}_{+}))$ $=\log\det(X(\mathbf{w}))+\log\left(\theta^{2}(\omega_{jk}^{2}-\omega_{k}\omega_{j})+\theta(\omega_{j}-\omega_{k})+1\right)$
If $\omega_{jk}^{2}>=\omega_{j}\omega_{k}$, we want to choose $\theta$ as big as possible while respecting the bounds. Otherwise, we want to choose $\theta$ such that the second $\log$ is maximized under consideration of the bound constraints.
$\theta^{*}$ =$=\begin{cases}\min\{u_{j}-w_{j},w_{k}-l_{k},\bar{\theta}\}&\text{if }\omega_{jk}^{2}<\omega_{j}\omega_{k}\\ \min\{u_{j}-w_{j},w_{k}-l_{k}\}&\text{if }\omega_{jk}^{2}\geq\omega_{k}\omega_{j}\end{cases}$
where $\bar{\theta}$ is defined as
$\bar{\theta}$ $=\frac{\omega_{j}-\omega_{k}}{2(\omega_{j}\omega_{k}-\omega_{jk}^{2})}.$
A-criterion
For the A-criterion, we get the following update in the objective.
@@ Appendix 0.B Notes on Custom Branch-and-Bound Algorithm for the Exact Optimal Design Problem
$D$ $=\omega_{j}\omega_{k}-\omega_{jk}^{2}.$
The objective value should be improving.

### v6

@@ A.2 Convexity of the objectives
By assumption, $p$ is positive as is the denominator. For the nominator, we use the Cauchy Schwartz inequality $\operatorname{Tr}\left(XY^{\intercal}\right)^{2}\leq\operatorname{Tr}(XX^{\intercal})\operatorname{Tr}(YY^{\intercal})$ and set $X=A^{-p/2}$ and $Y=A^{-\frac{p+2}{2}}B$
$h^{\prime\prime}(t)_{\restriction t=0}$ $\geq 0$
Thus, $k(X)$ is convex. ∎
A.3 Additional smoothness proofs
In following, we show that all the considered functions are $L$-smooth if we can bound the minimum eigenvalue of the input matrix from below.
Lemma A.25.
The function $f(X):=-\log\det(X)$ is $L$-smooth on $D:=\{X\in\mathbb{S}_{++}^{n}\,\mid\,\delta\leq\lambda_{\min}(X)\}$ with $L\geq\frac{1}{\delta^{2}}$.
Proof.
Let $A\mathbb{S}_{++}^{n}$, $B\in\mathbb{S}^{n}$ and $t\in[0,1]$. Then we can define $h(t)=-\log\det\left(A+tB\right)$. By Lemma A.22, $f(X)$ is convex and the second derivative of $h$ is given by
$h^{\prime\prime}(t)$ $=\sum\limits_{i=1}^{n}\frac{\lambda_{i}\left(A^{-1/2}BA^{-1/2}\right)^{2}}{1+t\lambda_{i}\left(A^{-1/2}BA^{-1/2}\right)}$
It is sufficient to show that $h^{\prime\prime}(t)\restriction_{t=0}\leq L$ to argue smoothness for $f$.
$h^{\prime\prime}(t)\restriction_{t=0}$ $=\sum\limits_{i=1}^{n}\lambda_{i}\left(A^{-1/2}BA^{-1/2}\right)^{2}$
By using a similar idea to Lemma A.29, we find
$\leq\sum\limits_{i=1}^{n}\lambda_{\max}\left(A^{-1}\right)^{2}\lambda_{i}(B)^{2}$
$=\lambda_{\max}\left(A^{-1}\right)^{2}\operatorname{Tr}\left(B^{2}\right)$
$=\lambda_{\max}\left(A^{-1}\right)^{2}\left\|B\right\|_{F}^{2}$
W.l.o.g. let $\left\|B\right\|_{F}=1$.
$=\frac{1}{\lambda_{\min}(A)^{2}}$
$\geq\frac{1}{\delta^{2}}$
$=L$
Thus, the function $f$ is smooth at any point $A\in D$. ∎
Lemma A.26.
For $p\in{\mathbb{R}}_{>0}$, the function $g(X):=\operatorname{Tr}\left(X^{-p}\right)$ is $L$-smooth on the domain $D:=\{X\in\mathbb{S}_{++}^{n}\,\mid\,\delta\leq\lambda_{\min}(X)\}$ with
$L\geq\frac{p(p+1)}{\delta^{p+2}}.$
Proof.
Let $A\in D$ and $B\in\mathbb{S}^{n}$ and $t\in[0,1]$. Define $h(t)=\operatorname{Tr}((A+tB)^{-p})$. In Lemma A.23 the proof for convexity of $g(X)$ is presented. There one can also find the calculation for $h^{\prime\prime}(t)_{\restriction t=0}$.
Since $A^{-p}$ is positive definite by Corollary A.19 and the trace is cyclic, we have
$=p(p+1)\operatorname{Tr}\left(A^{-\frac{p+2}{2}}BBA^{-\frac{p+2}{2}}\right)$
Note that $BB$ is postive semi-definite. Thus, we can use the bound in Fang et al. (1994).
$\leq p(p+1)\lambda_{\max}\left(A^{-(p+2)}\right)\operatorname{Tr}(BB)$
$=p(p+1)\lambda_{\max}\left(A^{-(p+2)}\right)\left\|B\right\|_{F}^{2}$
W.l.o.g. let $\left\|B\right\|_{F}=1$.
$=\frac{p(p+1)}{\lambda_{\min}\left(A^{p+2}\right)}$
$\geq\frac{p(p+1)}{\delta^{p+2}}$
Thus, $g$ is smooth with $L=\frac{p(p+1)}{\delta^{p+2}}$. ∎
Lemma A.27.
The function $k(X)=\log\left(\operatorname{Tr}(X^{-p})\right)$ is smooth on $D:=\{X\in\mathbb{S}_{++}^{n}\,\mid\,\alpha I\succcurlyeq X\succcurlyeq\delta I\}$.
Proof.
Let $A\in\mathbb{S}_{++}^{n}$, $B\in\mathbb{S}^{n}$ and $t\in{\mathbb{R}}$ s.t. $A+tB\in\mathbb{S}_{++}^{n}$. We define $h(t)=k(A+tB)=\log\left(\operatorname{Tr}((A+tB)^{-p})\right)$. By Lemma A.24, we have
$h^{\prime\prime}(t)_{\restriction t=0}$ $=p\frac{(p+1)\operatorname{Tr}\left(A^{-p-2}BB\right)\operatorname{Tr}\left(A^{-p}\right)-p\operatorname{Tr}\left(A^{-p-1}B\right)^{2}}{\operatorname{Tr}\left(A^{-p}\right)^{2}}$
$\leq p\frac{(p+1)\operatorname{Tr}\left(A^{-p}\right)\operatorname{Tr}\left(A^{-p-2}BB\right)}{\operatorname{Tr}(A^{-p})^{2}}$
$=p(p+1)\frac{\sum\limits_{i=1}^{n}\lambda_{i}\left(A^{-p-2}BB\right)}{\operatorname{Tr}\left(A^{-p}\right)}$
Note that $BB$ is PSD. The proof of Lemma A.29 can be adapted for one of matrices being symmetric and the other positive definite.
$\leq p(p+1)\frac{\lambda_{\max}\left(A^{-p-2}\right)\operatorname{Tr}(BB)}{\operatorname{Tr}\left(A^{-p}\right)}$
Observe that $Tr(BB)=\left\|B\right\|_{F}^{2}$. Without loss of generality, we can assume $\left\|B\right\|_{F}=1$.
$\leq p(p+1)\frac{\lambda_{\max}\left(A^{-p-2}\right)}{n\lambda_{\min}(A^{-p})}$
$=\frac{p(p+1)}{n}\frac{\lambda_{\max}(A^{p})}{\lambda_{\min}(A^{p+2})}$
$\leq\frac{p(p+1)}{n}\frac{\alpha^{p}}{\delta^{p+2}}$
$=L$
Remark.
In our case, the minimum eigenvalue is bounded around the optimum. However, it is not bounded in general for our feasible region unless we can introduce domain cuts. Hence, Lemmas A.27, A.25 and A.25 are not straight away useful to us.
A.4 Results on positive semi-definite matrices
We present two known results on the eigenvalues of the sum and product of positive definite matrices for completeness.
Lemma A.28 (Courant-Fischer min-max Theorem).
Let $A,B\in{\mathbb{R}}^{n\times n}$ symmetric matrices and let $\lambda_{1}(A)\leq\dots\leq\lambda_{n}(A)$ and $\lambda_{1}(B)\leq\dots\leq\lambda_{n}(B)$ denote their respective eigenvalues in increasing order. Then, we have the following inequalities concerning the eigenvalues of their sum $A+B$:
$\lambda_{k}(A)+\lambda_{1}(B)\leq\lambda_{k}(A+B)\leq\lambda_{k}(A)+\lambda_{n}(B)$ (12)
Proof.
Firstly, we know that
$\lambda_{1}(A)$ $\leq\frac{\left\langle A,\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}\leq\lambda_{n}(A)\quad\forall\mathbf{x}\in{\mathbb{R}}^{n}\backslash\{0\}.$
By Conrad-Fisher Min-Max Theorem, we have
$\lambda_{k}(A+B)$ $=\min_{\begin{subarray}{c}F\subset{\mathbb{R}}^{n}\\ \dim F=k\end{subarray}}\left(\max_{\mathbf{x}\in F\backslash\{0\}}\frac{\left\langle(A+B)\mathbf{x},\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}\right)$
$=\min_{\begin{subarray}{c}F\subset{\mathbb{R}}^{n}\\ \dim F=k\end{subarray}}\left(\max_{\mathbf{x}\in F\backslash\{0\}}\frac{\left\langle A\mathbf{x},\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}+\frac{\left\langle B\mathbf{x},\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}\right).$
Most likely, there will not be an $\mathbf{x}$ attaining the maximum simultaneously for summands. Using the above inequality for the eigenvalues, we get
$\geq\lambda_{1}(B)+\min_{\begin{subarray}{c}F\subset{\mathbb{R}}^{n}\\ \dim F=k\end{subarray}}\left(\max_{\mathbf{x}\in F\backslash\{0\}}\frac{\left\langle A\mathbf{x},\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}\right)$
$=\lambda_{1}(B)+\lambda_{k}(A).$
The proof in the other direction follows analogously. ∎
Lemma A.29.
Let $A,B\in\mathbb{S}_{++}^{n}$ and let $\lambda_{1}(A)\leq\dots\leq\lambda_{n}(A)$ and $\lambda_{1}(B)\leq\dots\leq\lambda_{n}(B)$ denote their respective eigenvalues in increasing order. Then, we have the following inequalities concerning the eigenvalues of $AB$:
$\lambda_{k}(A)\lambda_{1}(B)\leq\lambda_{k}(AB)\leq\lambda_{k}(A)\lambda_{n}(B)\quad\forall k\in[n].$ (13)
Proof.
Since $A$ and $B$ are both positive definite, they both have an invertible root, i.e. there exist matrices $\sqrt{A}$ and $\sqrt{B}$ with $A=\sqrt{A}\sqrt{A}$ and $B=\sqrt{B}\sqrt{B}$. Observe that $\lambda_{k}(AB)=\lambda_{k}(\sqrt{B}A\sqrt{B})$ and by definition of eigenvalues, we have
$\lambda_{1}(B)\leq\frac{\left\langle B\mathbf{x},\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}\leq\lambda_{n}(B)\quad\forall\mathbf{x}\in{\mathbb{R}}^{n}\backslash\{0\}$
. Thus, we find
$\lambda_{k}(AB)$ $=\lambda_{k}(\sqrt{B}A\sqrt{B})$
$=\min_{\begin{subarray}{c}F\subset{\mathbb{R}}^{n}\\ \dim F=k\end{subarray}}\left(\max_{\mathbf{x}\in F\backslash\{0\}}\frac{\left\langle\sqrt{B}A\sqrt{B}\mathbf{x},\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}\right)$
$=\min_{\begin{subarray}{c}F\subset{\mathbb{R}}^{n}\\ \dim F=k\end{subarray}}\left(\max_{\mathbf{x}\in F\backslash\{0\}}\frac{\left\langle A\sqrt{B}\mathbf{x},\sqrt{B}\mathbf{x}\right\rangle}{\left\langle\sqrt{B}\mathbf{x},\sqrt{B}\mathbf{x}\right\rangle}\frac{\left\langle B\mathbf{x},\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}\right)$
Using the above inequality for the eigenvalues of $B$ and the fact that the root of $B$ is invertible, we get
$\geq\lambda_{1}(B)\min_{\begin{subarray}{c}F\subset{\mathbb{R}}^{n}\\ \dim F=k\end{subarray}}\left(\max_{\mathbf{x}\in F\backslash\{0\}}\frac{\left\langle A\mathbf{x},\mathbf{x}\right\rangle}{\left\langle\mathbf{x},\mathbf{x}\right\rangle}\right)$
$=\lambda_{1}(B)\lambda_{k}(A)$
The other side of the statement follows analogously. Note that we technically only use that $B$ is regular. The matrix $A$ can be positive semi-definite. ∎
A.5 Notes on the Frobenius norm
Lemma A.30.
Let $A\in{\mathbb{R}}^{m\times n}$ and $B\in{\mathbb{R}}^{n\times p}$. Then,
$\sigma_{\min}(A)\|B\|_{F}\leq\|AB\|_{F}\leq\sigma_{\max}\|B\|_{F}$ (14)
Proof.
By the definition of the Frobenius norm, we have
$\|AB\|_{F}^{2}$ $=\operatorname{Tr}\left((AB)^{\intercal}AB\right)$
$=\operatorname{Tr}(A^{\intercal}AB^{\intercal}B).$
Note that both $A^{\intercal}A$ and $B^{\intercal}B$ are positve semi-definite. Thus, we can use the result in Fang et al. (1994):
$\geq\lambda_{\min}(A^{\intercal}A)\operatorname{Tr}(B^{\intercal}B)$
$=\sigma_{\min}(A)\|B\|_{F}^{2}.$
The norm is nonnegative in any case and by definition singular values are also nonnegative. Thus, we can take the root and get
$\sigma_{\min}(A)\|B\|_{F}\leq\|AB\|_{F}.$
The other direction follows analogously. ∎
B Custom branch-and-bound with coordinate descent for OEDP
In this section, we showcase the changes we made to the solver set-up in Ahipaşaoğlu (2021). Firstly, we state and solve the problems of interest as minimization problems. Secondly, we forgo the $\mathbf{w}N\in{\mathbb{Z}}\,\forall i\in[m]$ in favor of directly having the variables $x$ to encode if and how often the experiments should be run. The D-Optimal Problem is then stated as
$\min_{\mathbf{x}}\,$ $\,-\log\det\left((X(\mathbf{x}))\right)$
s.t. $\sum\limits_{i=1}^{m}x_{i}=N$
$\mathbf{x}\geq 0$
$x_{i}\in{\mathbb{Z}}\,\forall i\in[m].$
The A-Optimal Problem is formulated as
$\min_{\mathbf{x}}\,$ $\,\log\left(\operatorname{Tr}(X^{-1})\right)$
s.t. $\sum\limits_{i=1}^{m}x_{i}=N$
$\mathbf{x}\geq 0$
$x_{i}\in{\mathbb{Z}}\,\forall i\in[m].$
Additionally, we have adapted the step size for the A-Optimal Problem. For the update in the objective, we get
@@ B Custom branch-and-bound with coordinate descent for OEDP
$D$ $=\omega_{j}\omega_{k}-\omega_{jk}^{2}$
where $\omega_{j}=\mathbf{v}_{j}^{\intercal}X(x)^{-1}\mathbf{v}_{j}$ and $\zeta_{j}=\mathbf{v}_{j}^{\intercal}X(x)^{-2}\mathbf{v}_{j}$, and $\omega_{jk}=\mathbf{v}_{j}^{\intercal}X(x)^{-1}\mathbf{v}_{k}$ and $\zeta_{jk}=\mathbf{v}_{j}^{\intercal}X^{-2}\mathbf{v}_{k}$. The objective value should be improving.